ok please help this one is hard,
The difference in the age of two girls is 1 year. The sum of their ages is 27. What are their ages?

Answer by jb007mi6(53)

0K...one girl is x and the other is y
you know the following
{x-y=1
{x+y=27
so add both equations and you get 2x=28 so x=14
14-1=y
y=13
the two girls are 13 and 14